OpenStudy (anonymous):

Can you help me with this one? Essay: Show all work. A gardener wants to create a triangular garden in the shape of a right triangle with shortest side length x – 2y ft. and the middle length side x + 6y ft. What is an algebraic expression for the area of the garden? Be sure to multiply this out and express in simplest correct mathematical form, including units.

OpenStudy (anonymous):

A = 1/2 * b * h, and they give you base and height, so: A = 1/2 (x - 2y ft)(x + 6y ft) = 1/2(x^2 + 4xy - 12y^2) ft^2 = 1/2x^2 + 2xy - 6y^2 ft^2